OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a red marble, do not replace it, and pick another red marble?

OpenStudy (anonymous):

How many red marbles are there @4everjames ?

OpenStudy (anonymous):

@waterineyes 3

OpenStudy (anonymous):

Similar to last question, just some changes: In the beginning, the probability of a red marble is 3/10 After you've picked the first one and do not replace it, the probability of a red marble is 2/9. So your statement would be: Probability of red and red = Probability of 1st red * probability of 2nd red = (3/10) * (2/9) = 6/90 = 1/15

OpenStudy (anonymous):

Yes.. And what are total number of marbles that bag has?

OpenStudy (anonymous):

@waterineyes 10

OpenStudy (anonymous):

@mubzz thanksz!!!! i get it now thanks to u and @waterineyes

OpenStudy (anonymous):

An easy tip to solve all probability questions: When it says (prob of x) AND (prob of y) you multiply, i.e. (prob of x)*(prob of y) When it says (prob of x) OR (prob of y) you add, i.e. (prob of x)+(prob of y)

OpenStudy (anonymous):

ok just for review (A bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a red marble, replacing it, and then picking a black marble?